### v1

# ForAug: Recombining Foregrounds and Backgrounds to Improve Vision Transformer Training with Bias Mitigation

Abstract
Transformers, particularly Vision Transformers (ViTs), have achieved state-of-the-art performance in large-scale image classification. However, they often require large amounts of data and can exhibit biases that limit their robustness and generalizability. This paper introduces ForAug, a novel data augmentation scheme that addresses these challenges and explicitly includes inductive biases, which commonly are part of the neural network architecture, into the training data. ForAug is constructed by using pretrained foundation models to separate and recombine foreground objects with different backgrounds, enabling fine-grained control over image composition during training. It thus increases the data diversity and effective number of training samples. We demonstrate that training on ForNet, the application of ForAug to ImageNet, significantly improves the accuracy of ViTs and other architectures by up to 4.5 percentage points (p.p.) on ImageNet and 7.3 p.p. on downstream tasks. Importantly, ForAug enables novel ways of analyzing model behavior and quantifying biases. Namely, we introduce metrics for background robustness, foreground focus, center bias, and size bias and show that training on ForNet substantially reduces these biases compared to training on ImageNet. In summary, ForAug provides a valuable tool for analyzing and mitigating biases, enabling the development of more robust and reliable computer vision models. Our code and dataset are publicly available at https://github.com/tobna/ForAug.
1 Introduction
Figure 1: Comparison of ForNet and ImageNet. ForNet recombines foreground objects with different backgrounds each epoch, thus creating a more diverse training set. We still apply traditional data augmentation afterwards.
Image classification, a fundamental task in computer vision (CV), involves assigning a label to an image from a predefined set of categories. This seemingly simple task underpins a wide range of applications, including medical diagnosis [39, 50], autonomous driving [52], and object recognition [3, 17, 15]. Furthermore, image classification is used for large-scale pretraining of vision models [10, 31, 47] and to judge the progress of the field of CV [22, 36]. The advent of large-scale datasets, particularly ImageNet [8], containing millions of labeled images across thousands of categories, has been instrumental in driving significant progress in this field. ImageNet served as a catalyst for the rise of large-scale CV models [25, 16] and remains the most important CV benchmark for more than a decade [25, 48, 54, 16].
While traditionally, convolutional neural networks (CNNs) have been the go-to architecture for image classification, Transformers [49], particularly the Vision Transformer (ViT) [10], have emerged as a powerful alternative. These attention-based models have demonstrated superior performance in various vision tasks, including image classification [54, 58, 3, 63, 51].
Data augmentation is a key technique for training image classification models. Traditional data augmentation methods, such as random cropping, flipping, and color jittering, are commonly employed to increase the diversity of the training data and improve the model’s performance [56, 42]. These basic transformations, originally designed for CNNs, change the input images in a way that preserves their semantic meaning [1]. However, the architectural differences of CNNs and Transformers suggest that the latter might benefit from different data augmentation strategies. In particular, the Transformers self-attention mechanism is not translation equivariant [38, 9], meaning that the model does not inherently understand the spatial relationships between pixels.
Inspired by this inductive bias of CNNs, that is not inherent to ViTs, we propose ForAug, a novel data augmentation scheme for image classification which makes the translation equivariance of CNNs explicit in the training data by recombining foreground objects at varying positions with different backgrounds. Applying ForAug to ImageNet gives rise to ForNet, a novel dataset that enables this data augmentation with with fine-grained control over the image composition. Recognizing that Transformers need to learn the spatial relationships from data, since they are not inherently translation invariant, and in general are usually trained on larger datasets [24], we separate the foreground objects in ImageNet from their backgrounds, using an open-world object detector [37], and fill in the background in a plausible way using an object removal model [43, 45]. This allows us to recombine any foreground object with any background on the fly, creating a highly diverse training set. During recombination, we can control important parameters, like the size and position of the foreground object, to help the model learn the spatial invariances necessary for image classification. We show that training on ForNet instead of ImageNet increases the model accuracy of Transformers by up to 4.5 p.p. on ImageNet and an up to $39.3\%$ reduction in error rate on downstream tasks.
Additionally, ForAug is a useful tool for analyzing model behavior and biases, when used during the evaluation phase. We utilize our control over the image distribution to quantify a model’s background robustness (by varying the choice of background), foreground focus (by leveraging our knowledge about the placement of the foreground object), center bias (by controlling the object’s position), and size bias (by controlling object size). These analyses provide insights into model behavior and biases, which is crucial for model deployment and future robustness optimizations. We show that training on ForNet, instead of ImageNet, significantly reduces all of these biases, completely removing the models’ dependence on the background distribution. We make our code for ForAug and the ForNet-dataset publicly available [^1] https://github.com/tobna/ForAug to facilitate further research.
Contributions
  • We propose ForAug, a novel data augmentation scheme, that recombines objects and backgrounds to train Transformers for image classification.
  • We show that training on ForNet, the ImageNet instantiation of ForAug, leads to 4.5 p.p. improved accuracy on ImageNet and 7.3 p.p. on downstream tasks.
  • We propose novel ForAug-based metrics to analyze and quantify fine-grained biases trained models: Background Robustness, Foreground Focus, Center Bias, and Size Bias. Training on ForNet, instead of ImageNet, significantly reduces these biases.
2 Related Work
Data Augmentation for Image Classification
Data augmentation is a crucial technique for improving the performance and generalization of image classification models. Traditional augmentation strategies rely on simple geometric or color-space transformations like cropping, flipping, roatation, blurring, color jittering, or random erasing [62] to increase the diversity of the training data without changing their semantic meaning. With the advent of Transformers, new data augmentation operations like PatchDropout [30] have been proposed. Other transformations like Mixup [61], CutMix [59], or random cropping and patching [46] combine multiple input images. These simple transformations are usually bundled to form more complex augmentation policies like AutoAugment [5] and RandAugment [6], which automatically search for optimal augmentation policies or 3-augment [48] which is optimized to train a ViT. For a general overview of data augmentation techniques for image classification, we refer to [42, 56].
We build upon these general augmentation techniques by introducing a novel approach to explicitly separate and recombine foregrounds and backgrounds for image classification. Our approach is used in tandem with traditional data augmentation techniques to improve model performance and reduce biases.
Copy-Paste Augmentation
The copy-paste augmentation [14], which is used for object detection [41, 14] and instance segmentation [53, 28], involves copying segmented objects from one image and pasting them onto another. While typically human-annotated segmentation masks are used to extract the foreground objects, other foregound sources have been explored, like 3D models [19] and pretrained object-detection models for use on objects on white background [11] or synthetic images [12]. DeePaste [53] focuses on using inpainting for a more seamless integration of the pasted object.
Unlike these methods, ForNet focuses on image classification. While for detection and segmentation, objects are pasted onto another image (with a different foreground) or on available or rendered background images of the target scene, we extract foreground objects and fill in the resulting holes in the background in a semantically neutral way. This way, we can recombine any foreground object with a large variety of neutral backgrounds from natural images, enabling a controlled and diverse manipulation of image composition.
Model robustness evaluation
Evaluating model robustness to various image variations is critical for understanding and improving model generalization. Datasets like ImageNet-C [18] and ImageNet-P [18] introduce common corruptions and perturbations. ImageNet-E [27] evaluates model robustness against a collection of distribution shifts. Other datasets, such as ImageNet-D [60], focus on varying background, texture, and material, but rely on synthetic data. Stylized ImageNet [13] investigates the impact of texture changes. ImageNet-9 [55] explores background variations using segmented images, but the backgrounds are often artificial.
In contrast to these existing datasets, which are used only for evaluation, ForNet provides fine-grained control over foreground object placement, size, and background selection, enabling a precise and comprehensive analysis of specific model biases within the context of a large-scale, real-world image distribution. As ForNet also provides controllable training set generation, it goes beyond simply measuring robustness to actively improving it through training.
3 RecombiNet (Method)
Figure 2: Overview of ForNet. The data creation consists of two stages: (1, offline) Segmentation, where we segment the foreground objects from the background and fill in the background. (2, online) Recombination, where we combine the foreground objects with different backgrounds to create new samples.
We introduce ForAug, a data augmentation scheme designed to enhance Transformer training by explicitly separating and recombining foreground objects and backgrounds. ForAug involves two stages: Segmentation and Recombination, both visualized in Figure 2.
Segmentation
The segmentation stage isolates the foreground objects and their corresponding backgrounds. We then fill in the background in a visually plausible way [43] using a pretrained object-removal model. This stage is computed once offline and the results are stored for the recombination stage.
First, foreground objects are detected and segmented from their backgrounds using a prompt-based segmentation model to exploit the classification datasets labels. We use the state-of-the-art Grounded SAM [37], which is based on Grounding DINO [29] and SAM [23]. The prompt we use is “a <class name>, a type of <object category>”, where <class name> is the specific name of the objects class as defined by the dataset and <object category> is a the broader category of the object. The <object category> guides the segmentation model towards the correct object in case the <class name> alone is too specific. This can be the case with prompts like “sorrel” or “guenon”, where the more general name “horse” or “monkey” is more helpful. We derive the <object category> from the WordNet hierarchy, using the immediate hypernym.
We iteratively extract up to $n$ foreground masks for each dataset-image, using different more and more general prompts based on the more general synsets of WordNet (e.g. “a sorrel, a type of horse”, “a horse, a type of equine”, …). Masks that are very similar, with a pairwise IoU of at least $0.9$, are merged. The output is a set of masks delineating the foreground objects and the backgrounds. We select the best mask per image (according to Equation 1) in a later filtering step, described below.
An inpainting model that is specifically optimized to remove objects from images, such as LaMa [45] or Attentive Eraser [43], is used to inpaint the foreground regions in the backgrounds. To ensure the quality of the foreground and background images (for each dataset-image), we select a foreground/background pair from the $\leq n$ variants we have extracted and infilled in the previous steps. Using an ensemble of six ViT, ResNet, and Swin Transformer models pretrained on the original dataset, we select the foreground/background pair that maximizes foreground performance while minimizing the performance on the background and size of the foreground according to:
$\begin{split}\text{score}(\mathrm{fg},\mathrm{bg},c)&=\log\left(\frac{1}{\left|E\right|}\sum_{m\in E}\mathbb{P}[m(\mathrm{fg})=c]\right)\\ &+\log\left(1-\frac{1}{\left|E\right|}\sum_{m\in E}\mathbb{P}[m(\mathrm{bg})=c]\right)\\ &+\lambda\log\left(1-\left|\frac{\operatorname{size}(\mathrm{fg})}{\operatorname{size}(\mathrm{bg})}-\varepsilon\right|\right).\end{split}$ (1)
Here, $E$ is the ensemble of models and $m$ is a pretrained model, $c$ is the correct foreground class, $\mathrm{fg}$, and $\mathrm{bg}$ are the foreground and background and $\operatorname{size}(\cdot)$ is the size in number of pixels. We ran a hyperparameter search using a manually annotated subset of foreground/background variants to find the factors in Equation 1: $\lambda=2$ and $\varepsilon=0.1$. The optimal foreground size of $10\%$ of the full image balances the smallest possible foreground size that encompasses all the respective class information in the image with still conveying the foreground information after pasting it onto another background. This filtering step ensures we segment all the relevant foreground objects.
Finally, we filter out backgrounds that are more than $80\%$ infilled, as these tend to be overly synthetic, plain and don’t carry much information (see Appendix C). We ablate this choice in Section 4.1. In summary, we factorize the dataset into a set of foreground objects with a transparent background and a set of diverse backgrounds per class. The next step is to recombine them as data augmentation before applying common data augmentation operations during training.
Recombination
The recombination stage, which is performed online, combines the foreground objects with different backgrounds to create new training samples. For each object, we follow the pipeline of: Pick an appropriate background, resize it to a fitting size, place it in the background image, smooth the transition edge, and apply other data augmentations.
For each foreground object, we sample a background using one of the following strategies: (1) the original image background, (2) the set of backgrounds from the same class, or (3) the set of all possible backgrounds. These sets are trading off the amount of information the model can learn from the background against the diversity of new images created. In each epoch, each foreground object is seen exactly once, but a background may appear multiple times.
The selected foreground is resized based on its relative size within its original image and the relative size of the original foreground in the selected background image. The final size is randomly selected from a 30% range around upper and lower limits ($s_{u}$ and $s_{l}$), based on the original sizes:
$s\sim\mathcal{U}\left[(1-0.3)s_{l},(1+0.3)s_{u}\right].$ (2)
To balance the size of the foreground and that of the backgrounds original foreground, the upper and lower limit $s_{u}$ and $s_{l}$ are set to the mean or range of both sizes, depending on the foreground size strategy: mean or range.
The resized foreground is then placed at a random position within the background image. This position is sampled from a generalization of the Bates distribution [2] with parameter $\eta\in\mathbb{N}$, visualized in Figure 3. We choose the bates distribution, as it presents an easy way to sample from a bounded domain with just one hyperparameter that controls the concentration of the distribution. $\eta=1$ corresponds to the uniform distribution; $\eta>1$ concentrates the distribution around the center; and for $\eta<-1$, the distribution is concentrated at the borders. To more seamlessly integrate the foreground, we apply a Gaussian blur with ${\sigma\in[\frac{\sigma_{\text{max}}}{10},\sigma_{\text{max}}]}$, inspired by the standard range for the Gaussian blur operation in [48], to the foreground’s alpha-mask.
We can apply standard data augmentation techniques in two modes: Either we apply all augmentations to the recombined image, or we apply the cropping and resizing to the background only and then apply the other augmentations after recombination. The second mode ensures the foreground object remains fully visible, while the first mode mirrors standard data augmentation practices.
We experiment with a constant mixing ratio, or a linear or cosine annealing schedule that increases the amount of images from the original dataset over time. The mixing ratio acts as a probability of selecting an image from the original dataset; otherwise, an image with the same foreground is recombined using ForAug. Thus, we still ensure each foreground is seen once per epoch.
4 Experiments
We conduct a comprehensive suit of experiments to validate the effectiveness of our approach. We compare training on ForNet, the ImageNet instantiation of ForAug, to training on ImageNet for 7 different models. Furthermore, we assess the impact of using ForNet for pretraining on multiple fine-grained downstream datasets. Additionally, we use ForAug’s control over the image distribution to quantify some model behaviors and biases.
4.1 Design Choices of ForAug
We start by ablating the design choices of ForAug. For this, we revert to TinyImageNet [26], a subset of ImageNet containing 200 categories with 500 images each, and TinyForNet, a version of ForAug derived from TinyImageNet. Table 1 presents the results of these ablations.
Dataset Detect. Infill FG. Augmentation BG. BG. edge original TinyImageNet Accuracy
prompt Model size Order strategy pruning smoothing image mixing ViT-Ti [%] ViT-S [%]
TinyImageNet $66.1\pm 0.5$ $68.3\pm 0.7$
[…]
Table 1: Ablation of design decisions of TinyForNet on TinyImageNet and ForNet on ImageNet.
Prompt. First, we evaluate the type of prompt used to detect the foreground object. Here, the general prompt, which contains the class and the more general object category, outperforms only having the class name (specific).
Inpainting. Attentive Eraser [43] produces superior results compared to LaMa [45] (see Appendix B for examples).
Foreground size significantly impacts performance. Employing a range of sizes during recombination, rather than a fixed mean size, boosts accuracy by approximately 1 p.p. This suggests that the added variability is beneficial.
[…]
Background pruning. When it comes to the choice of backgrounds to use, we test two pruning thresholds ($t_{\text{prune}}$) to exclude backgrounds with excessive inpainting. A threshold of $t_{\text{prune}}=1.0$ means that we use all backgrounds that are not fully infilled. Varying $t_{\text{prune}}$ has minimal impact. Therefore, we choose $t_{\text{prune}}=0.8$ to exclude predominantly artificial backgrounds. Similarly, applying edge smoothing to foreground masks with Gaussian blurring actually hurts performance on TinyForNet, but slightly improves it on ForNet.
Mixing ForNet with the original ImageNet data proves crucial. While constant and linear mixing schedules improve performance over no mixing by $2-3$ p.p. compared to only using TinyForNet, the cosine annealing schedule yields the best results, boosting accuracy by another $0.5-1$ p.p.
Background strategy. Another point is the allowed choice of background image for each foreground object. We compare using the original background, a background from the same class, and any background. These strategies go from low diversity and high shared information content between the foreground and background to high diversity and low shared information content. For ViT-Ti, the latter two strategies perform comparably, while ViT-S benefits from the added diversity of using any background. The same is true when training on the full (ImageNet) version of ForNet.
Figure 3: Plot of the probability distribution function (PDF) of the extended Bates distribution for different parameters $\eta$. Higher values of $\eta$ concentrate the distribution around the center.
Training Set/ Bates Parameter TIN TinyForNet
$\eta=-3$ $-2$ $/-1$ 1 $2$ $3$
TinyImageNet 68.9 60.5 60.2 60.8 62.6 63.1
$\eta=-3$ 71.3 79.3 79.5 79.1 79.3 79.1
$\eta=-2$ 71.5 80.0 78.7 79.3 79.1 78.8
$\eta=1/-1$ 72.3 79.5 78.9 80.2 79.7 80.4
$\eta=2$ 71.3 78.2 77.8 79.1 79.6 79.9
$\eta=3$ 71.4 77.2 76.9 78.6 79.6 79.7
Table 2: Accuracy of ViT-S trained on TinyImageNet (TIN) and TinyForNet with different foreground position distributions by varying the parameter of a Bates distribution $\eta$. The best performance is achieved using the uniform distribution ($\eta=1$).
Foreground position. Finally, we analyze the foreground object’s positioning in the image. We utilize an extended Bates distribution to sample the position of the foreground object. The Bates distribution [2] with parameter $\eta\geq 1$ is the mean of $\eta$ independent uniformly distributed random variables [20]. Therefore, the larger $\eta$, the more concentrated the distribution is around the center. We extend this concept to $\eta\leq-1$ by defining ${X\sim\text{Bates}(\eta):\Leftrightarrow s(X)\sim\text{Bates}(-\eta)}$ for $\eta\leq 1$ with $s$ being the sawtooth function on $[0,1]$:
$s(x)=\begin{cases}x+0.5&\text{if }0<x<0.5\\ x-0.5&\text{if }0.5\leq x\leq 1\end{cases}$ (3)
Note that $s\circ s=\mathrm{id}$ on $[0,1]$. This way, distributions with $\eta\leq-1$ are more concentrated around the borders. $\eta=1$ and $\eta=-1$ both correspond to the uniform distribution. The PDF of this extended Bates distribution is visualized in Figure 3.
When sampling more towards the center of the image, the difficulty of the task is reduced, which then reduces the performance on TinyImageNet. This is reflected in the performance when evaluating on TinyForNet with $\eta=2$ and $\eta=3$ compared to $\eta=-1/1$. We observe a similar reduction for $\eta<-1$. This experiment is conducted using the LaMa infill model.
Dataset Classes Training Images Validation Images
TinyImageNet 200 100,000 10,000
TinyForNet 200 99,404 9,915
ImageNet 1,000 1,281,167 50,000
ForNet 1,000 1,274,557 49,751
Table 3: Dataset statistics for TinyImageNet, TinyForNet, ImageNet, and ForNet. For ForNet and TinyForNet we report the number of foreground/background pairs.
After fixing the optimal design parameters in Table 1 (last row), we construct the full ForNet dataset using the entire ImageNet dataset. Table 3 compares the dataset statistics of ImageNet and ForNet. The slightly reduced image count in ForNet is due to instances where Grounded SAM failed to produce valid object detections.
4.2 Image Classification Results
Model ImageNet Accuracy when trained on Delta
ImageNet ForNet
ViT-S $79.1\pm 0.1$ $81.4\pm 0.1$ +2.3
ViT-B $77.6\pm 0.2$ $81.1\pm 0.4$ +3.5
ViT-L $75.3\pm 0.4$ $79.8\pm 0.1$ +4.5
Swin-Ti $77.9\pm 0.2$ $79.7\pm 0.1$ +1.8
Swin-S $79.4\pm 0.1$ $80.6\pm 0.1$ +1.2
ResNet-50 $78.3\pm 0.1$ $78.8\pm 0.1$ +0.5
ResNet-101 $79.4\pm 0.1$ $80.4\pm 0.1$ +1.0
Table 4: ImageNet results of models trained on ForNet and on ImageNet directly. ForNet improves the performance of all models in our test.
Table 4 compares the ImageNet performance of models trained on ForNet and ones trained directly on ImageNet. We adopt the training setup of [33] and [48] (details in Appendix A) for training ViT [10], Swin [31] and ResNet [16] models. Notably, ForNet improves performance across all tested architectures, including the ResNet models (up to $1$ p.p.), demonstrating benefits beyond Transformers. For Transformer models, we observe improvements from $1.2$ p.p. to $4.5$ p.p. This improvement is more substantial for the larger models, with ViT-L gaining $4.5$ p.p. in accuracy. ForNet’s improvements mostly counteract the drop in performance due to overfitting for large models. When training on ImageNet, this drop is $3.8$ p.p. from ViT-S to ViT-L, while for ForNet it is reduced to $1.6$ p.p.
Model Aircraft Cars Flowers Food Pets
ViT-S @ ImageNet $72.4\pm 1.0$ $89.8\pm 0.3$ $94.5\pm 0.2$ $89.1\pm 0.1$ $93.8\pm 0.2$
ViT-S @ ForNet $78.6\pm 0.5$ $92.2\pm 0.2$ $95.5\pm 0.2$ $89.6\pm 0.1$ $94.5\pm 0.2$
+6.2 +2.4 +1.0 +0.5 +0.7
ViT-B @ ImageNet $71.7\pm 0.5$ $90.0\pm 0.2$ $94.8\pm 0.4$ $89.8\pm 0.2$ $94.1\pm 0.4$
ViT-B @ ForNet $79.0\pm 2.2$ $93.3\pm 0.1$ $96.5\pm 0.1$ $90.9\pm 0.1$ $95.1\pm 0.4$
+7.3 +3.3 +1.7 +1.1 +1.0
ViT-L @ ImageNet $72.1\pm 1.0$ $88.8\pm 0.3$ $94.4\pm 0.3$ $90.1\pm 0.2$ $94.2\pm 0.4$
ViT-L @ ForNet $77.6\pm 1.2$ $89.1\pm 0.2$ $96.6\pm 0.1$ $91.3\pm 0.1$ $95.1\pm 0.1$
+5.5 +0.3 +2.2 +1.2 +0.9
Swin-Ti @ ImageNet $77.0\pm 0.1$ $91.3\pm 0.6$ $95.9\pm 0.1$ $90.0\pm 0.2$ $94.2\pm 0.1$
Swin-Ti @ ForNet $81.1\pm 0.8$ $92.8\pm 0.4$ $96.2\pm 0.1$ $90.4\pm 0.3$ $94.8\pm 0.5$
+4.1 +2.5 +0.3 +0.4 +0.6
Swin-S @ ImageNet $75.7\pm 1.4$ $91.0\pm 0.3$ $95.9\pm 0.5$ $91.1\pm 0.2$ $94.4\pm 0.1$
Swin-S @ ForNet $81.4\pm 0.2$ $93.1\pm 0.2$ $96.3\pm 0.3$ $91.2\pm 0.2$ $94.9\pm 0.3$
+5.7 +2.1 +1.4 +0.1 +0.5
ResNet-50 @ ImageNet $78.2\pm 0.5$ $89.8\pm 0.2$ $91.7\pm 0.4$ $84.4\pm 0.2$ $93.7\pm 0.3$
ResNet-50 @ ForNet $80.3\pm 0.4$ $90.4\pm 0.2$ $91.7\pm 0.2$ $84.5\pm 0.2$ $93.7\pm 0.3$
+2.1 +0.6 $\pm$0 +0.1 $\pm$0
ResNet-101 @ ImageNet $78.4\pm 0.6$ $90.3\pm 0.1$ $91.2\pm 0.5$ $86.0\pm 0.2$ $94.3\pm 0.2$
ResNet-101 @ ForNet $81.4\pm 0.5$ $91.3\pm 0.1$ $92.9\pm 0.2$ $86.3\pm 0.1$ $94.0\pm 0.3$
+3.0 +1.3 +1.7 +0.3 -0.3
Table 5: Downstream accuracy in percent when finetuning on other datasets. Models were pretrained on ForNet and ImageNet. Pretraining on ForNet increases Transformer downstream accuracy on all datasets.
To assess the transferability of ForNet-trained models, we finetune models pretrained on ImageNet and ForNet on five fine-grained datasets: FGVC-Aircraft [32], Stanford Cars [7], Oxford Flowers [34], Food-101 [21], and Oxford-IIIT Pets [35]. While for ResNets, the performance of both training datasets is about the same, for every Transformer, we see the accuracy improve on all downstream dataset by up to 7.3 p.p. and a reduction of error rate of up to $39.3\%$. In summary, these results demonstrate that the improved representation learning achieved by training on ForNet translates to superior performance not only on ImageNet, but also on a variety of fine-grained image classification tasks.
4.3 Further Model Evaluation
Beyond its use for training, ForNet’s unique properties and controlled data generation capabilities make it a powerful tool for analyzing model behavior and biases.
Background Robustness
Model Background Robustness when trained on Delta
ImageNet ForNet
ViT-S $0.73\pm 0.01$ $0.99\pm 0.01$ +0.26
ViT-B $0.72\pm 0.01$ $1.00\pm 0.01$ +0.28
ViT-L $0.70\pm 0.01$ $1.00\pm 0.01$ +0.30
Swin-Ti $0.72\pm 0.01$ $1.00\pm 0.01$ +0.28
Swin-S $0.72\pm 0.01$ $1.00\pm 0.01$ +0.28
ResNet-50 $0.79\pm 0.01$ $0.99\pm 0.01$ +0.20
ResNet-101 $0.79\pm 0.01$ $1.00\pm 0.01$ +0.21
Table 6: Evaluation of the background robustness of models trained on ForNet and on ImageNet directly. Training on ForNet improves the background robustness of all model to $\approx 1.00$, meaning the model is indifferent to the choice of background.
We assess the robustness of models to shifts in the background distribution from a class-related background to any background. Background robustness is defined to be the ratio of accuracy on ForNet with same-class backgrounds to accuracy with any background:
Background $\text{Background Robustness}=\frac{\text{Acc}(\textit{ForNet}_{\text{all}})}{\text{Acc}(\textit{ForNet}_{\text{same}})}$ (4)
It represents the relative drop in performance under a background distribution shift. Table 6 presents the background robustness of various models. When trained on ImageNet, smaller models generally exhibit greater robustness to changes in the background distribution than larger models and ResNet is more robust than the tested Transformer models. Crucially, training on ForNet instead of ImageNet improves the background robustness of all models to $\approx 1.00$, meaning that these models are agnostic to the choice of background and only classify based on the foreground. These findings highlight the generalization benefits of ForNet.
Foreground Focus
Model Foreground Focus when trained on
IN FN IN FN IN FN
GradCam GradCam++ IG
ViT-S $1.2\pm 0.1$ $2.3\pm 0.3$ $1.2\pm 0.1$ $2.1\pm 0.4$ $1.9\pm 0.1$ $2.7\pm 0.1$
ViT-B $1.2\pm 0.1$ $2.4\pm 0.7$ $1.1\pm 0.1$ $2.1\pm 0.1$ $1.7\pm 0.1$ $2.7\pm 0.1$
ViT-L $1.3\pm 0.1$ $1.6\pm 0.1$ $1.1\pm 0.1$ $1.3\pm 0.1$ $1.3\pm 0.1$ $2.6\pm 0.1$
Swin-Ti $0.9\pm 0.1$ $0.7\pm 0.1$ $1.0\pm 0.3$ $0.7\pm 0.3$ $2.5\pm 01$ $4.8\pm 0.3$
Swin-S $0.8\pm 0.1$ $0.7\pm 0.1$ $0.7\pm 0.1$ $0.7\pm 0.4$ $2.4\pm 0.1$ $4.6\pm 0.3$
ResNet-50 $2.2\pm 0.1$ $2.7\pm 0.1$ $2.0\pm 0.1$ $2.9\pm 0.1$ $3.2\pm 0.1$ $4.9\pm 0.2$
ResNet-101 $2.3\pm 0.1$ $2.8\pm 0.1$ $2.2\pm 0.1$ $3.0\pm 0.1$ $3.2\pm 0.1$ $4.8\pm 0.1$
Table 7: Evaluation of the foreground focus using GradCam, GradCam++ and IntegratedGradients of models trained on ForNet (FN) and on ImageNet (IN) directly. Training on ForNet improves the foreground focus of almost all models.
Leveraging our inherent knowledge of the foreground masks when using ForNet, as well as common XAI techniques [40, 4, 44], we can evaluate a model’s focus on the foreground object. We can directly evaluate ImageNet trained models, but this technique can also be extended to other datasets without relying on manually annotated foreground-masks. To evaluate the foreground focus, we employ Grad-CAM [40], Grad-CAM++ [4] or IntegratedGradients (IG) [44] to compute the per-pixel importance of an image for the model’s prediction. The foreground focus is defined to be the ratio of the foreground’s relative importance to its relative size in the image:
[…]
The foreground focus of a model is its average foreground focus over all test images. Table 7 presents our findings. Training on ForNet significantly increasees the foreground focus of ViT and ResNet across all metrics used. For Swin, the foreground focus stagnates when measured using GradCam and GradCam++, but almost doubles when using IG.
Center Bias
Model Center Bias when trained on Delta
ImageNet ForNet
ViT-S [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$0.255\pm 0.008$ $0.220\pm 003$0.035
ViT-B [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$0.254\pm 0.004$ $0.190\pm 0.002$0.064
ViT-L [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$0.243\pm 0.011$ $0.117\pm 0.007$0.126
Swin-Ti [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$0.250\pm 0.007$ $0.165\pm 0.002$0.085
Swin-S [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$0.232\pm 0.001$ $0.156\pm 002$0.076
ResNet50 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$0.263\pm 0.003$ $0.197\pm 0.003$0.066
ResNet101 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$0.230\pm 0.003$ $0.199\pm 002$0.031
Table 8: Evaluation of the position bias. We plot the accuracy relative to the center accuracy of multiple instantiations of the models when the foreground objects is in different cells a $3\times 3$ grid. Training on ForNet significantly reduces a models center bias.
With ForNet we have unique control over the position of the foreground object in the image. This lets us quantify the center bias of ImageNet- and ForNet-trained models. We divide the image into a $3\times 3$ grid and evaluate model accuracy when the foreground object is in each of the $9$ grid cells. Each cell’s accuracy is divided by the accuracy in the center cell for normalization, which gives us the relative performance drop when the foreground is in each part of the image. The center bias is calculated as one minus the average of the minimum performance of a corner cell and the minimum performance of a side cell:
Center $\begin{split}&\text{Center Bias}=\\ &\hskip 7.0pt1-\frac{\min\limits_{a,b\in\{0,2\}}\text{Acc}(\text{cell}_{(a,b)})+\min\limits_{\begin{subarray}{c}a=1\text{ or }b=1\\ a\neq b\end{subarray}}\text{Acc}(\text{cell}_{(a,b)})}{2\text{Acc}(\text{cell}_{(1,1)})}\end{split}$ (6)
Table 8 visualizes the center bias of three instantiations of each model. Performance is generally highest in the center and the center top and bottom and center left and right cells, and lowest in the four corners. Interestingly, ImageNet-trained models perform slightly better when the foreground object is on the right side of the image, compared to the left side, despite our use of random flipping with a probability of $0.5$ during training. Training on ForNet significantly reduces center bias across all models. This demonstrates that ForNet promotes a more uniform spatial attention distribution. Their accuracy is higher in the center left and right cells than in the center top and bottom ones, which is not the case for ImageNet-trained models.
Size Bias
Figure 4: Evaluation of the size bias of models trained on ForNet. We plot the accuracy relative to the accuracy when using the mean foreground size.
Finally, we evaluate the impact of different-sized foreground objects on the accuracy. For this evaluation, we use the mean foreground size strategy. We introduce a size factor $f_{\text{size}}$ by which we additionally scale the foreground object before pasting it onto the background. Results are again normalized by the accuracy when using the mean foreground size ($f_{\text{size}}=1.0$). Figure 4 shows the size bias curves of ViT-S and ViT-B when trained on ImageNet and ForNet. Models trained on ForNet maintain better performance even with smaller foreground objects, when ImageNet-trained models exhibit a more rapid performance decline. Therefore, ForNet-training improves robustness to variations in object scale.
5 Discussion & Conclusion
We introduce ForAug, a novel data augmentation scheme that facilitates improved Transformer training for image classification. By explicitly separating and recombining foreground objects and backgrounds, ForAug enables controlled data augmentation, leading to significant performance gains on ImageNet and downstream fine-grained classification tasks. Furthermore, ForAug provides a powerful framework for analyzing model behavior and quantifying biases, including background robustness, foreground focus, center bias, and size bias. Our experiments demonstrate that training on ForNet, the instantiation of ForAug on ImageNet, not only boosts accuracy but also significantly reduces these biases, resulting in more robust and generalizable models. In the future, we see ForAug be also applied to other datasets and tasks, like video recognition or segmentation. ForAug’s ability to both improve performance and provide insights into model behavior makes it a valuable tool for advancing CV research and developing more reliable AI systems.
Acknowledgements
This work was funded by the Carl-Zeiss Foundation under the Sustainable Embedded AI project (P2021-02-009) and by the EU project SustainML (Horizon Europe grant agreement No 101070408). All compute was done thanks to the Pegasus cluster at DFKI.
Appendix A Training Setup
Parameter Value
Image Resolution $224\times 224$
Epochs 300
Learning Rate 3e-3
Learning Rate Schedule cosine decay
Batch Size 2048
Warmup Schedule linear
Warmup Epochs 3
Weight Decay 0.02
Label Smoothing 0.1
Optimizer Lamb [57]
Data Augmentation Policy 3-Augment [48]
Table 9: Training setup for our ImageNet and ForNet training.
Dataset Batch Size Epochs Learning Rate
Aircraft 512 500 3e-4
Cars 1024 500 3e-4
Flowers 256 500 3e-4
Food 2048 100 3e-4
Pets 512 500 3e-4
Table 10: Training setup for finetuning on different downstream datasets. Other settings are the same as in Table 9.
On ImageNet we use the same training setup as [33] and [48] without pretraining. As our focus is on evaluating the changes in accuracy due to ForAug/ForNet, like [33], we stick to one set of hyperparameters for all models. We list the settings used for training on ImageNet and ForNet in Table 9 and the ones used for finetuning those weights on the downstream datasets in Table 10.
Appendix B Infill Model Comparison
LaMa Att. Eraser LaMa Att. Eraser
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 11: Example infills of LaMa and Attentive Eraser.
Appendix C Images with High Infill Ratio
Infill Ratio LaMa Att. Eraser
93.7 [Uncaptioned image] [Uncaptioned image]
95.7 [Uncaptioned image] [Uncaptioned image]
83.7 [Uncaptioned image] [Uncaptioned image]
88.2 [Uncaptioned image] [Uncaptioned image]
Table 12: Example infills with a large relative foreground area size that is infilled (infill ratio).